three shirts and 2 pants cost $50. two shirts and 3 pants cost $80. what is the cost of one shirt and one pants?
Answer by mananth(16946)

three shirts and 2 pants cost $50.
two shirts and 3 pants cost $80.
..
let shirt cost $x & pants cost $y
..
3x+2y=50..............1
2x+3y=80..............2
..
multiply equation 1 by -2 & equation 2 by 3 to eliminate x
..
-2(3x+2y)+3(2x+3y)= -2*50+3*80
-6x-4y+6x+9y=-100+240
+5y=140
y=$28 cost of pants
plug value of y in equation 1
3x+2y=50
3x+2*28=50
3x+56=50
3x=-6
x=-2 .
X cannot be negative .
However cost of 1 Pant and 1 shirt = $26 ( mathematically)